solve by the elimination method 0.05x +0.25y=44 0.15x +0.05y=48
Answer by stanbon(75887)

solve by the elimination method
0.05x +0.25y=44
0.15x +0.05y=48
----
1st get rid of the decimals:
5x + 25y = 4400
15x + 5y = 4800
------
2nd: Multiply thru the 1st equation by 3:
15x + 75y = 3*4400
15x + 5y = 4800
------
Subtract and solve for "y":
70y = 8400
y = 120
-----
Solve for "x":
5x + 25y = 4400
5x + 25*120 = 4400
5x + 3000 = 4400
5x = 1400
x = 280
-------------------
Cheers,
Stan H.